Winnie the Pooh is going to visit the Rabbit, who lives a kilometer away from him. He walked there at a speed of v = 2 km / h, and, having been a guest for a while, well-fed and satisfied, returned back at a speed D, = 1 km / h. How long did Winnie the Pooh spend visiting the Rabbit if the average speed of his way from home and back is v = 1 km / h?
1
Expert's answer
2020-12-21T11:02:43-0500
"v=\frac{2d}{t_1+t_2+t}\\v=\frac{2d}{\frac{d}{v_1}+\frac{d}{v_2}+t}
\\1=\frac{2(1)}{\frac{1}{2}+\frac{1}{1}+t}\\t=0.5\ h"